The sum of two numbers is 72. One of the numbers is six more than twice the other. What are the two numbers?
Found 2 solutions by gsmani_iyer, atif.muhammad:
Answer by gsmani_iyer(201)


Let the smaller number is = x

So the greater number = 2x+6

The sum of the numbers = x+2x+6 = 72

So 3x+6 = 72

3x = 72-6 = 66

So x = 66/3 = 22;

2x+6 = 2*(22)+6 = 50

The smaller number = 22; the greater number = 50. Answer

Answer by atif.muhammad(135)

The sum of two numbers is 72.  One of the numbers is six more than twice the other.  What are the two numbers?


Let one number be x, and the other be y.

The sum of two numbers is 72   ------   x+y = 72

One of the numbers is six more than twice the other.  ---   x = 2y + 6

We have a pair of simultaneous equations, which we can solve.

x+y = 72....(1)
x = 2y + 6....(2)

Substitute (2) into (1):

(2y + 6) + y = 72

3y + 6 = 72

3y = 66

y = 22...(3)

Subsitute (3) into (2):

x = 2(22) + 6 = 50

Hence, our two numbers are 50,22.